[Halld-jlab] Hall D Staff Meeting Notes, July 28, 2014
Mark M. Ito
marki at jlab.org
Mon Jul 28 19:09:09 EDT 2014
People,
Find the notes below and at
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_28,_2014
.
-- Mark
______________________________________________________________________
Present: Eugene Chudakov, Mark Dalton, Alexandre Deur, Hovanes Egiyan,
Sergey Furletov, Mark Ito, David Lawrence, Lubomir Pentchev, Yi Qiang,
Elton Smith (chair), Steve Smith, Sascha Somov, Simon Taylor, Glenn
Young, Beni Zihlmann
=Announcements (Elton)=
* PAC is going on now
** tomorrow talks from Mike Williams on DIRC, Liping Gan on rare eta decays
* Readiness Review report is final, but we have not seen it yet
** likely we will be asked to transition draft documents to final form
* Aug. 8: photon beam commissioning review
** Chair: Stepan Stepanyan, committee members: Mike Spata, Brian[?] Freeman
** preliminary agenda exists
** Patrizia Rossi wants more on rad con issues given tagger hall
experience of last April
*** Pawel Degtiarenko will give presentation
* Aug. 26-28: Accelerator Readiness Review
** conducted by DOE, outside reviewers
** last Jan. Hall A had to
presenthttps://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_28,_2014
** this time Hall B (HPS) and Hall D (GlueX)
* Hovanes on schedule:
** HPS in the Fall
** Proton Charge Radius in the Spring
=Safety Culture Survey (Steve)=
* results of the June '14 cycle presented
* have gone to a 6 month cycle from 3 months
* 23 of 86 possible respondents from Physics
* work planning and control an issue for Physics
* areas for possible improvement:
** task lists
** worker involvement in planning
** willingness to stop work
* [full presentation will be linked from this page]
=DAQ and Online (Sergey F. and David)=
* continuing search for source of hanging gluon notes
* all tests show a hardware problem
* individual tests (memory, CPU, etc.) succeed
* problem only seen with complex configurations
* no crashes seen with gluon53, gluon01, gluon02a
* new memory coming in today or tomorrow, will see if that makes difference
* CODA is in common with all failures seen so far
* direct DMA from network to memory a suspect
* Elton suggested mapping out location of standard online application on
screens in control room
** David and Hovanes will make a proposal
* code developed for pulse pedestal object
** they are generated by firmware but have not been available before
** Mark D. noted that average of four samples truncated back to
one-sample precision
=Trigger (Sascha)=
* config files: F1TDCs done last week
** used south side LED pulser
** no FIFO overflows, chip/channel correspondence correct
** F1 chip status show errors
** parity reported, not in data stream
** data format has redundancies, clean up later
* BCAL trigger had wrong rate
** due to blown fuse in TI board
** William Gu diagnosed it
** master TI replaced, one slave board as well
** problem ween in same crate before
** can give config files to Sergey F. now
* will now turn to global trigger work
* plan to move FDC boards to original place
=Drift Chambers (Beni)=
* took data over weekend, no crashes
* looking at data now, lots of things to do
* database loaded with proper DAC values
* noise
** CDC especially, but FDC too
** hundreds of ADC counts if noise
** CAEN HV supplies likely cause
** Mike Staib looking at CDC
=CDC Alignment (Simon)=
* success in using data to do alignment
=FDC (Lubomir)=
* analyzing data, mostly vertical cosmics
** several wires hit per plane, strip data not very useful (too many strips)
** will do a check of go or no go for each channel
** comparing previous results, see some channels go from bad to good and
some from good to bad within the overall 0.5% problem rate
* very few 25 package events
** got two from Beni's horizontal trigger rig
** very useful in identifying problems
* getting 600 micron resolution
* efficiency very high
* Eugene: magnet will be cold soon, need a plan for detector tests with
field on
=Offline (Mark)=
* all makefile for sim-recon have been deleted
* started working on data challenge 3
* re-started work on global build of GlueX software to include version
configuration
* Richard Jones proceeding with changes to mcsmear
* Richard Jones starting work on "smearing" global time in HDGeant
=Beamline (Alexandre)=
* pair spectrometer field monitoring
** monitor current, temp. of coil
** difficult to install NMR
** will do test to see if this is enough
* draft of run plan written
** shown to Elton
** three positions for gamma profiler planned: before collimator, before
magnet, before beam dump
** next step: detailed procedures for tasks in run plan
* material for new radiators arrived
* plan in place for radiation monitoring for tagger hall and Hall D
* commissioning target
** four targets, two of them plastic, one carbon, one of varying thickness
** gamma profiler can be used as an extreme upstream target
=Slow Controls (Hovanes)=
* Nerses and Vanik rolling out changes to HV and LV applications
* chillers for BCAL and FCAL: original plan PLC, instead Wesley Moore
working on EPICS application
* Manuel Lara working on FCAL voltage control in EPICS
* converter stage in FEL
** Yi and H. testing
** then to alignment group, stick position and motor setting correlation
* radiator ladder surveyed and installed
** targets to be installed when ready
** controls simple, code in place
* collimator controls installed
* TAC: new motor ordered, due beginning of August
* Sascha: active collimator
** trying to put missing wedges
** in touch with Tom Carstens
** design is tricky
** Alex Barnes back at UConn
** crate and controller in TEDF
** they should take care of everything including power supply
=Solenoid (Yi and Eugene)=
* magnet will be cold in a week or two
* then execute commissioning plan
** stress magnet for a week or two
** test detectors
** will do a fast dump
* need to rope off perimeter
* need to guard against flying metal objects
=Beam Schedule (Eugene)=
* nominal date: October 10
* 3 weeks total running planned, not including breaks
* will need break to install equipment in tagger hall after initial push
of beam into hall
* next beam: likely March
--
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295
More information about the Halld-jlab
mailing list